OpenStudy (anonymous):

If the geometric mean of two +ve numbers a and b is 10, then logz+logb=?? what does geometric mean mean? @phi @ganeshie8

OpenStudy (phi):

geometric mean comes from a,b,c having a common ratio \[ \frac{a}{b} = \frac{b}{c} \] if you cross multiply you get \[ b^2 = a c \\ b= \sqrt{ac}\] in your case you have \[ \sqrt{a b} = 10 \]

OpenStudy (phi):

in other words, you have a G.P. a, 10, b with a common ratio of a/10 (or 10/b)

OpenStudy (anonymous):

what does geometric mean mean?...

OpenStudy (phi):

First, it means what I posted up above. It probably comes from similar right triangles (this is an aside)